OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (campbell_st):

this is a quadratic equation \[7\sin^2(x) - 14\sin(x) + 7 = 0\] which can simplify to \[7(\sin^2(x) - 2\sin(x) + 1) = 0\] the quadratic can be factorised and is a perfect square. so that should make it easier to find the solution.

OpenStudy (anonymous):

I'm really stuck on this question can you elaborate more? :))

OpenStudy (campbell_st):

well make a simple substitution let sin(x) = u so your problem is \[7(u^2 -2u +1) = 0\] all you need to do is factor the quadratic in the brackets...

OpenStudy (anonymous):

So i get (u-1)(u-1)

OpenStudy (anonymous):

I'm really confused please help